Find the volume of the solid generated by revolving the circle x^2 + y^2 = 4 about the line x = 3. Use the cylindrical shell method.


Note: With sketch please
1
Expert's answer
2020-05-06T19:28:34-0400

Consider "x^2+y^2=4, y\geq0." Then "y=\sqrt{4-x^2}"

Figure shows the region and a cylindrical shell formed by rotation about the line "x=3." It has radius "3-x," circumference "2\pi(3-x)," and height "\sqrt{4-x^2}."


The volume of the given solid is


"V_1=\displaystyle\int_{-2}^22\pi(3-x)\sqrt{4-x^2}dx"

"x=2\sin{t},\ -\pi/2\leq t\leq \pi/2"

"\ dx=2\cos{t}dt"

"\sqrt{4-x^2}=2\cos{t}"


"\int2\pi(3-x)\sqrt{4-x^2}dx="

"=2\pi \int(3-2\sin{t})(2\cos{t})(2\cos{t})dt="

"=24\pi \int\cos^2{t}dt-16\pi\int\sin{t}\cos^2{t}dt="

"=12\pi \int(1+\cos{2t})dt+16\pi ({\cos^3{t} \over 3})="

"=12\pi t+6\pi \sin{2t}+{16\pi \over 3}\cos^3{t}+C="

"=12\pi\arcsin{({x \over 2})}+3\pi x\sqrt{4-x^2}+{2\pi \over 3}(4-x^2)^{3/2}+C"

"V_1=\pi\bigg[12\arcsin{({x \over 2})}+3x\sqrt{4-x^2}+{2\over 3}(4-x^2)^{3/2}\bigg]\begin{matrix}
 2 \\
 -2
\end{matrix}="

"=\pi(12({\pi \over 2})+0+0-12(-{\pi \over 2})-0-0)=12\pi^2 (units^3)"

"V=2V_1=2(12\pi^2)=24\pi^2(units^3)"

The volume of the solid generated by revolving the circle "x^2+y^2=4" about the line "x=3" is "24\pi^2" cubic units.